OpenStudy (anonymous):

Medal!! Help me please Solve. 3(4x − 2x) = 8

OpenStudy (tkhunny):

Have you considered resolving the parentheses?

OpenStudy (anonymous):

You mean like, 3x4?

OpenStudy (tkhunny):

No, I mean like 4 - 2 = 2

OpenStudy (anonymous):

12x-6x=8 6x=8 x=3/4

OpenStudy (tkhunny):

You had a 4 and you then have a 12? Why would you want to go that way? Seriously, resolve the parentheses.

OpenStudy (tkhunny):

In the beginning, the worst number we have is an 8. Can we keep it that way? 3(4x-2x) = 8 Resolve the parentheses 3(2x) = 8 Divide by 2 3x = 4

OpenStudy (anonymous):

Is the answer x= 4/3? ..

OpenStudy (anonymous):

yes, sorry i had that backward on mine

OpenStudy (anonymous):

Is that correct @tkhunny

OpenStudy (anonymous):

Nvm...

OpenStudy (tkhunny):

You tell me. Work with confidence.